Find any extrema for: f(x)= -x^-5x+3 Is this maximum or a minimum?
This is how far I got, is it correct?
x= -[x^-5x+6.25] 6.25+3
x= -[x-2.5]^ + 9.25
x=2.5 y=9.25 This is a maximum

Answer by edjones(8007)

No.
Although your final answer is correct your work shows that you don't understand how to get there.
f(x)=y=-x^2-5x+3 I assume that this is the problem. You didn't put a number for the power in the equation you submitted.
If the coefficient of x^2 is negative then it is a maximum and vise versa.
The formula for the maxima or minima at x is x=-b/2a
x=-(-5)/(2*-1)
=5/-2
=-5/2=-2.5
y=-2.5^2-(5*-2.5)+3
=-6.25+12.5+3
=9.25
.
Ed